# Low-Rank Online Dynamic Assortment with Dual Contextual Information

### v1

@@ 4.2 Main Results
We first establish an estimation error bound on our rank-constrained maximum likelihood estimation for the LOAD problem.
@@
Suppose Assumptions 1, 2 and 3 hold. Further assume we have an initial estimate $\Phi_{0}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ such that $\|\Phi_{0}-\Phi^{*}\|_{F}\leq c_{2}\sqrt{\sigma_{r}^{*}}$ for some constant $c_{2}$. Let $\hat{\Phi}$ be the rank-constrained estimator using Algorithm 1 with initial estimate $\Phi_{0}$. If the exploration length $T_{0}$ satisfies
@@
for some constant $C_{1},C_{2}$ and any $\delta\in(0,1)$, $B>0$, where $\Sigma=\Sigma_{q}\otimes\Sigma_{p}$, then
@@
Note that we require the initialization error $\|\Phi_{0}-\Phi^{*}\|_{F}^{2}$ to be bounded. This assumption, often referred to as the ‘basin of attraction’ is a standard practice in the non-convex low-rank matrix estimation literature (Jain et al., , 2013; Chi et al., , 2019). It ensures that the iterative algorithm converges to the desired target. Such assumptions are valid in diverse situations, such as when offline data are available, or when the length of the exploration stage is large enough to provide a warm start for the gradient descent algorithm.
@@
Suppose Assumptions 1, 2 and 3 hold. Further assume that $\|\Theta^{*}_{r+1:d_{2},r+1:d_{1}}\|_{F}^{2}\leq S_{\perp}$ for some constant $S_{\perp}$. Let $\widehat{\theta}_{n,rtv}$ be the maximum likelihood estimator of $\theta^{*}_{rtv}$ using observations from user $t=1,\ldots,n$ by minimizing (5). Then,
$\|\widehat{\theta}_{n,rtv}-\theta^{*}_{rtv}\|_{W_{n}}$ $\leq\frac{1}{\kappa}\left(\sigma\sqrt{2df\log\left(1+\frac{n}{df}\right)+4\log{n}}+S_{2}^{2}\cdot S_{\perp}\cdot\sqrt{nK}\right)=:\alpha_{n}$
with probability at least $1-1/n^{2}$, where $W_{n}=\sum_{t=1}^{n}\sum_{i\in S_{t}}x_{it,rtv}x_{it,rtv}^{\top}$ and $df=(d_{1}+d_{2})r-r^{2}$ is the dimension of the reduced space.
@@
Theorem 1 provides a detailed regret bound of our policy. When applying our policy to e-commerce applications with a large number of users (large $T$), we can further simply this rate. With the choice $B=b\sqrt{T}$ for some constant $b$, we have $S_{\perp}=O(r/\sqrt{T})$ by definition and therefore $\operatorname*{\mathcal{R}}_{T}=\tilde{O}(r(d_{1}+d_{2})\sqrt{T})$, where $\tilde{O}(\cdot)$ ignores some logarithm term. For $d$-dimensional contextual assortment problem, Chen et al., (2020) established an $\Omega(d\sqrt{T}/K)$ lower bound on the cumulative regret. As the capacity $K$ is the maximum number of items that are displayed each time, we can interpret $K$ as a constant, and consequently, the lower bound becomes $\Omega(d\sqrt{T})$. When these results are applied by interpreting vectorization of the outer product of dual features as a single feature, the lower bound becomes $\Omega(d_{1}d_{2}\sqrt{T})$, without the low-rank assumption. Our algorithm significantly reduces the rate of regret to $\tilde{O}(df\cdot\sqrt{T})$ in the low-rank regime.
Establishing this regret bound possesses two main technical challenges. The first challenge involves incorporating the low-rank structure. To address this, we utilize results from the low-rank matrix estimation literature and adapt them to the LOAD problem. This enables us to establish an estimation bound on the estimated subspace. The second challenge arises from the inherent combinatorial structure of the assortment problem. To overcome this, we construct bound on the difference between the reward in the original problem and the transformed problem in the reduced space, which enables us to establish regret bounds for our algorithm.
4.3 Proof Sketch of Regret Bound
This section provides an outline of the three major steps in the proof of Theorem 1. Detailed proofs are presented in the Supplementary Materials. Step 1, as shown in Proposition 1, establishes a bound on the bias induced from rotation and truncation of the parameter matrix. Step 2 shows a bound on the difference between the optimistic utility $z_{it}$ defined as (4) and the true utility $v_{it}$ (Lemma 2). Finally, using this bound on the optimistic utility, Step 3 derives a bound on the regret of the $t$-th user, and the final result on the bound of cumulative regret (Theorem 1).
Proposition 2 (Bounds for Subspace Estimation).
$\|\Theta^{*}_{r+1:d_{2},r+1:d_{1}}\|_{F}^{2}\leq\frac{288\sigma_{1}^{3}}{\sigma_{r}^{3}\cdot\kappa}\cdot\left(\frac{3}{4}+\frac{8T_{0}}{\kappa B}+\frac{2}{K\|\Sigma\|_{2}}\right)\frac{Kr\log(\frac{d_{1}+d_{2}}{\delta})}{B}S_{2}^{4}=:S_{\perp}$ (6)
with probability at least $1-\delta$.
Lemma 2 (Bound on Optimistic Utility).
with probability at least $1-t^{-2}$.
$r_{t}=R(\{v_{it}\},S_{t}^{*})-R(\{v_{it}\},S_{t})\leq R(\{z_{it}\},S_{t}^{*})-R(\{v_{it}\},S_{t})\leq R(\{z_{it}\},S_{t})-R(\{v_{it}\},S_{t}).$ (7)
Moreover, we can bound the last term using the result from Oh and Iyengar, (2021) that establishes a bound on the difference of expected reward of an assortment given different utilities:
Lemma 3 (Lemma 5 of Oh and Iyengar, (2021)).
Combining (7), Lemmas 2 and 3, we have
${\operatorname*{\mathcal{R}}}_{T}$ $=\sum_{t=1}^{T_{0}}r_{t}+\sum_{t=T_{0}+1}^{T}r_{t}$
Lastly, presenting an upper bound on $\sum_{t=T_{0}+1}^{T}2\alpha_{t}\max_{i\in S_{t}}\|x_{it,rtv}\|_{W_{t}^{-1}}$, we obtain our final result Theorem 1. Proofs of the theoretical results are presented in the Supplementary Materials.
@@ 5 Numerical Studies
In this section, we evaluate the cumulative regret of our ELSA-UCB policy across varying scenarios, including different numbers of users, feature dimensions, ranks of parameter matrix, numbers of items and capacity of the assortment. We begin by examining it using synthetic data in Section 5.1 and subsequently validate its practicability in a real-world dataset in Section 5.2.
@@ 5.1 Simulation
To generate the parameter matrix $\Phi^{*}$, we first construct a diagonal matrix $D^{*}$ with $r$ non-zero components, each set to be 10. Next we multiply random orthonormal matrices $U^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times r}$ and $V^{*}\in\operatorname*{\mathbb{R}}^{d_{1}\times r}$ to generate a random rank-$r$ matrix $\Phi^{*}=U^{*}D^{*}(V^{*})^{\top}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$. For the user features, we generate $q_{t}\in\operatorname*{\mathbb{R}}^{d_{1}}$ for every $t\in[T]$ from the $d_{1}$-dimensional standard Gaussian distribution. Similarly, for item features, we generate $p_{i}\in\operatorname*{\mathbb{R}}^{d_{2}}$ from the $d_{2}$-dimensional standard Gaussian distribution for every $i\in[N]$. Note that the $N$ item features are fixed over the time horizon while user features are generated at each time step. The revenue is set as $r_{i}=1$ for every $i\in[N]$, following the setup in Oh and Iyengar, (2021). This accounts for scenarios where the goal is to maximize the sum of a binary outcome, aiming to maximize the total number of conversions, such as clicks or purchases in online retail.
Utilizing the generated parameters and features, we simulate user choices and the corresponding assortment selections following the policy. We compare the performance of our proposed ELSA-UCB against two versions of the UCB-MNL (Agrawal et al., , 2019; Oh and Iyengar, , 2021). As UCB-MNL considers features for each user-item pair as $x_{it}$, we consider two formulations of the joint features; “stacked” and “vectorized”.
@@
We apply our proposed ELSA-UCB policy and the two variations of UCB-MNL as baselines for comparison. We repeat the simulation 10 times and plot the median and the 10th/90th percentile of the cumulative regret for each $T\in\{500,1000,2000,5000,10000\}$ under different environments.
Figure 5: Cumulative regrets of our proposed ELSA-UCB algorithm compared with Stacked UCB-MNL and Vectorized UCB-MNL for different ranks $r\in\{1,5,10\}$.
For the first line of experiments, we fix $d_{1}=30,d_{2}=20,N=20,K=3$ and change the ranks $r\in\{1,5,10\}$ and number of users over $T=500,1000,2000,5000,10000$ and plot the corresponding cumulative regret in Figure 5. In higher ranks, the Stacked UCB-MNL produces larger cumulative regret, as it does not take consideration for the diverse interactions between the dual features. On the other hand, the Vectorized UCB-MNL has comparable performance when the rank is high. However, when the rank is low, the performance of Vectorized UCB-MNL is not satisfactory as it does not take account for the low-rank structure of the parameter matrix. Our ELSA-UCB policy universally outperforms both methods as it utilizes the low-rank structure and therefore efficiently estimate how the interactions between the dual features affect user preference of items.
@@
In the next set of experiments, we fix $d_{2}=20,r=3,N=20,K=3$ while changing $d_{1}\in\{30,50,100\}$. The results are shown in Figure 6. The Stacked UCB-MNL shows poor performance across all dimensions. Moreover, we can observe that the gap between the cumulative regret of ELSA-UCB and Vectorized UCB-MNL widens as $d_{1}$ increases. This indicates that our algorithm is even preferable in environments with higher dimensional features. Furthermore, note that the scale of the regret increases as $d_{1}$ increases which aligns with the results of Theorem 1.
@@
In the final set of experiments, we fix $d_{1}=30,d_{2}=20,r=3$ and change $(N,K)$ to be $\{(20,3),(100,3),(100,20)\}$. The results are illustrated in Figure 7. Note that as the assortment capacity $K$ increases, the gap between the optimal assortment and the near-optimal assortments decreases, resulting in smaller regrets. Overall, our proposed method outperforms both Stacked UCB-MNL and Vectorized UCB-MNL across all setups.
@@ 5.2 Real Data Application
In this section, we apply our algorithm to the “Expedia Hotel” dataset (Adam et al., , 2013) to evaluate its performance in real data.
@@ 5.2.1 Data Pre-processing
It is important to note that search queries constrain assortments to those within the search destination. We focus on data from the top destination with the highest search volume. Additionally, we filter out columns with missing over $90\%$ of searches. To address missing values in the remaining data for average star ratings and prices of each user from historical data, we employ a simple regression tree imputation. For hotel features, we compute the mean of feature values across searches. We discretize numerical features such as star ratings or review scores into categories to improve model fitting. After pre-processing, we have $T=4465$ unique searches encompassing $N=124$ different hotels, with $d_{2}=10$ hotel features and $d_{1}=18$ user features. The description of features is given in Table 1. We normalize each feature to have mean 00 and variance $1$. As for the item features, we truncate the feature values between $[-3,3]$ to avoid the effect of outliers.

### v2

@@ 4.2 Main Results
First we establish an estimation error bound on our rank-constrained maximum likelihood estimation for the LOAD problem.
@@
Suppose Assumptions 1, 2 and 3 hold. Further assume we have an initial estimate $\Phi_{0}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ such that $\|\Phi_{0}-\Phi^{*}\|_{F}\leq c_{2}\sqrt{\sigma_{r}}$ for some constant $c_{2}$, where $\sigma_{r}$ is $r$-th largest singular value of $\Phi^{*}$. Let $\hat{\Phi}$ be the rank-constrained estimator using Algorithm 1 with initial estimate $\Phi_{0}$. If the exploration length $T_{0}$ satisfies
@@
for some constants $C_{1},C_{2}$ and any $\delta\in(0,1)$, $B>0$, where $\Sigma=\Sigma_{q}\otimes\Sigma_{p}$, then
@@
Note that we require the initialization error $\|\Phi_{0}-\Phi^{*}\|_{F}^{2}$ to be bounded. This assumption, often referred to as the ‘basin of attraction’ is a standard practice in the non-convex low-rank matrix estimation literature (Jain et al., 2013; Chi et al., 2019). It ensures that the iterative algorithm converges to the desired target. Such assumptions are valid in diverse situations, such as when offline data are available, or when the length of the exploration stage is large enough to provide a warm start for the gradient descent algorithm.
The initial estimator $\Phi_{0}$ can also be chosen as the unconstrained maximum likelihood estimator (MLE) based on data collected during the exploration phase. Under standard regularity conditions for the MNL model, the unconstrained MLE achieves the convergence rate: $\|\hat{\Phi}_{MLE}-\Phi^{*}\|_{F}=O(\sqrt{d_{1}d_{2}/T_{0}})$ (Van der Vaart, 2000). Therefore, if the exploration length $T_{0}$ satisfies $T_{0}\gtrsim d_{1}d_{2}/\sigma_{r}^{2}$, the unconstrained MLE will satisfy the initial condition.
@@
Suppose Assumptions 1, 2 and 3 hold. Further assume that $\|\Theta^{*}_{r+1:d_{2},r+1:d_{1}}\|_{F}^{2}\leq S_{\perp}$ for some $S_{\perp}$. Let $\widehat{\theta}_{n,rtv}$ be the maximum likelihood estimator of $\theta^{*}_{rtv}$ using observations from user $t=1,\ldots,n$ by minimizing (6). Then,
$\|\widehat{\theta}_{n,rtv}-\theta^{*}_{rtv}\|_{W_{n}}$ $\leq\frac{1}{\kappa nK}\left(\sigma\sqrt{2df\log\left(1+\frac{n}{df}\right)+4\log{n}}+S_{2}^{2}\cdot S_{\perp}\cdot\sqrt{nK}\right)=:\alpha_{n}$
with probability at least $1-1/n^{2}$, where $W_{n}=\frac{1}{nK}\sum_{t=1}^{n}\sum_{i\in S_{t}}x_{it,rtv}x_{it,rtv}^{\top}$ and $df=(d_{1}+d_{2})r-r^{2}$ is the dimension of the reduced space.
@@
[…]
Establishing this regret bound possesses two main technical challenges. The first challenge involves incorporating the low-rank structure. To address this, we utilize results from the low-rank matrix estimation literature and adapt them to the LOAD problem. This enables us to establish an estimation bound on the estimated subspace. The second challenge arises from the inherent combinatorial structure of the assortment problem. To overcome this, we construct bound on the difference between the rewards in the original problem and the transformed problem in the reduced space, which enables us to establish regret bounds for our algorithm. We provide a proof sketch in Section S.7.1 of the Appendix.
@@ 5 Numerical Studies
In this section, we evaluate the cumulative regret of our ELSA-GIC policy across varying scenarios, including different numbers of users, feature dimensions, ranks of parameter matrix, numbers of items and capacity of the assortment. We begin by examining it using synthetic data in Section 5.1 and subsequently validate its practicability in a real-world dataset in Section 5.2.
@@ 5.1 Simulation
To generate the parameter matrix $\Phi^{*}$, we first construct a diagonal matrix $D^{*}$ with $r$ non-zero components, each set to be 10. Next we multiply random orthonormal matrices $U^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times r}$ and $V^{*}\in\operatorname*{\mathbb{R}}^{d_{1}\times r}$ to generate a random rank-$r$ matrix $\Phi^{*}=U^{*}D^{*}(V^{*})^{\top}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$. For the user features, we generate $q_{t}\in\operatorname*{\mathbb{R}}^{d_{1}}$ for every $t\in[T]$ from the $d_{1}$-dimensional standard Gaussian distribution. Similarly, for item features, we generate $p_{i}\in\operatorname*{\mathbb{R}}^{d_{2}}$ from the $d_{2}$-dimensional standard Gaussian distribution for every $i\in[N]$. Note that the $N$ item features are fixed over the time horizon while user features are generated at each time step. The revenue is set as $r_{i}=1$ for every $i\in[N]$, following the setup in Oh and Iyengar (2021). This accounts for scenarios where the goal is to maximize the sum of binary outcomes, aiming to maximize the total number of conversions, such as clicks or purchases in online retail.
Utilizing the generated parameters and features, we simulate user choices and the corresponding assortment selections following the policy. We compare the performance of our proposed ELSA-UCB with the estimated rank from GIC (denoted ELSA-GIC) against two versions of the UCB-MNL (Agrawal et al., 2019; Oh and Iyengar, 2021). As UCB-MNL considers features for each user-item pair as $x_{it}$, we consider two formulations of the joint features; “stacked” and “vectorized”.
@@
We apply our proposed ELSA-GIC policy and the two variations of UCB-MNL as baselines for comparison. We repeat the simulation 50 times and plot the mean and 95% confidence interval of the cumulative regret for each $T\in\{1000,2000,3000,5000,7500,10000\}$ under different environments.
First we demonstrate the consistency of rank-estimation through GIC (4). We fix $d_{1}=30,d_{2}=20,N=20,K=3,r=3$ and change the number of users over $T=100,200,500,1000,2000$ to compare the estimated rank when $T$ increases. As shown in Figure 5, the probability of accurate rank estimation nearly converges to $1$ as the sample size increases.
Figure 5: Accuracy of estimated rank using GIC with increasing $T$.
Figure 6: Cumulative regrets of our proposed ELSA-GIC algorithm compared with Stacked UCB-MNL and Vectorized UCB-MNL for different ranks $r\in\{2,5,10\}$.
For the first line of experiments, to assess the performance in terms of cumulative regret, we fix $d_{1}=50,d_{2}=20,N=20,K=3$ and change the rank $r\in\{2,5,10\}$ and number of users over $T=1000,2000,3000,5000,7500,10000$ and plot the corresponding cumulative regret in Figure 6. In higher ranks, the Stacked UCB-MNL produces larger cumulative regret, as it does not take consideration for the diverse interactions between the dual features. On the other hand, the Vectorized UCB-MNL has comparable performance when the rank is high. However, when the rank is low, the performance of Vectorized UCB-MNL is not satisfactory. Our ELSA-GIC policy universally outperforms both methods as it utilizes the low-rank structure and therefore efficiently estimate how the interactions between the dual features affect user preference of items. We also note that the Stacked method corresponds to a special case with $r=2$, which is only correctly specified when the true parameter matrix has nonzero entries solely in the main-effect terms, the first row and the first column. In our simulations, however, the parameter matrix includes nonzero interaction terms, so the Stacked method is mis-specified even when the true rank equals $2$. Our method, in contrast, remains correctly specified in this case and achieves superior performance.
@@
In the next set of experiments, we fix $d_{2}=20,r=3,N=20,K=3$ while changing $d_{1}\in\{30,50,100\}$. The results are shown in Figure 7. The Stacked UCB-MNL shows poor performance across all dimensions. Moreover, we can observe that the gap between the cumulative regret of ELSA-GIC and Vectorized UCB-MNL widens as $d_{1}$ increases. This indicates that our algorithm is even preferable in environments with higher dimensional features. Furthermore, note that the scale of the regret increases as $d_{1}$ increases which aligns with the results of Theorem 1.
@@
In the third set of experiments, we fix $d_{1}=30,d_{2}=20,r=3$ and change $(N,K)$ to be $\{(20,3),(100,3),(100,20)\}$. The results are illustrated in Figure 8. Note that as the assortment capacity $K$ increases, the gap between the optimal assortment and the near-optimal assortments decreases, resulting in smaller regrets. Overall, our proposed method outperforms both Stacked UCB-MNL and Vectorized UCB-MNL across all setups.
In the final set of experiments, we evaluate the robustness of ELSA-GIC under model misspecification. We consider three scenarios:
  1. 1.
    Approximately low-rank $\Phi^{*}$: The first $r$ singular values of $\Phi^{*}$ are fixed at a constant level, while the remaining singular values are set to be small - specifically, 1% of the leading singular values - capturing approximate low-rank structure with residual noise.
  2. 2.
    Full-rank $\Phi^{*}$: The parameter matrix is drawn as a full-rank matrix without dominant singular directions, favoring the Vectorized UCB-MNL model, which does not assume any low-rank structure.
  3. 3.
    Main-effect-only model: The utility is generated without any interaction between user and item features and is influenced solely by the main effects. In this setting, the effective rank is 2, and the structure favors the Stacked UCB-MNL model.
We fix $(d_{1},d_{2},N,K)=(30,20,20,3)$, and use $r=3$ for the approximately low-rank matrix. As shown in Figure 9, ELSA-GIC achieves the lowest regret in the approximately low-rank case, demonstrating its strength when the low-rank assumption only holds approximately. Moreover, it remains competitive even when the data-generating process favors alternative models, achieving comparable regret to the specialized methods under their respective advantageous regimes.
Figure 9: Cumulative regret over time $T$ under various model misspecification settings.
@@ 5.2 Real Data Application
In this section, we apply our algorithm to the “Expedia Hotel” dataset (Adam et al., 2013) to evaluate its performance in real data.
@@ 5.2.1 Data Pre-processing
It is important to note that search queries constrain assortments to those within the search destination. We focus on data from the top destination with the highest search volume. Additionally, we filter out columns with missing over $90\%$ of searches. To address missing values in the remaining data for average star ratings and prices of each user from historical data, we employ a simple regression tree imputation. For hotel features, we compute the mean of feature values across searches. We discretize numerical features such as star ratings or review scores into categories to improve model fitting. After pre-processing, we have $T=4465$ unique searches encompassing $N=124$ different hotels, with $d_{2}=10$ hotel features and $d_{1}=18$ user features. The description of features is given in Table 1. We normalize each feature to have mean 0 and variance $1$. As for the item features, we truncate the feature values to lie within [3,3][-3,3] to avoid the effect of outliers.